OpenStudy (anonymous):

A student receives his grade report from a local community college, but the GPA is smudged. He took the following classes: a 2 hour credit art, a 3 hour credit history, a 4 hour credit science course, a 3 hour credit mathematics course, and a 1 hour science lab. He received a “B” in the art class, an “A” in the history class, a “C” in the science class, a “B” in the mathematics class, and an “A” in the science lab. What was his GPA if the letter grades are based on a 4 point scale? (A=4, B=3, C=2, D=1, F=0)

OpenStudy (anonymous):

for each subject, multiply the scale of the grade by the credit hours and add them all up, then divide them by the total number of hours..

OpenStudy (anonymous):

that is: 2B+3A+4C+3B+1A=2(3)+3(4)+4(2)+3(3)+1(4)=39 points now divide this by the total hours which is 13 39/13=3 so the GPA is 3 out of 4

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

I understand now. Thanks for the help.

OpenStudy (anonymous):

you're welcome.. I wish you an A in your exam ;)